OpenStudy (anonymous):

Find the outside surface area. Note that the container may not have a top. Assume a = 2 in., b = 1 in., c = 1.4 in., d = 2 in., e = 1.5 in.

OpenStudy (anonymous):

excuse me, what do a, b, c, d, and e stand for?

OpenStudy (anonymous):

ya, what are a, b, c, d, and e

OpenStudy (anonymous):

and you did not mention the very thing is a sphere or cube or something else

OpenStudy (anonymous):

sorry im tryiing to get it right now

OpenStudy (anonymous):

The three rectangles: Area = ab + bc + bd The two triangles: Area = (1/2)ae + (1/2)ae = ae

OpenStudy (anonymous):

good, so you can see that the object is a triangular prism, specifically here with two identical triangles parallel to each other vertically. notice that the sides of the tri-prism are all rectangulars. set the outside surface area as S. then S=S1+S2+S3+S4(S1 stands for the area of the bottom triangle, S2, S3, S4 the areas of the front, left and the right sides.)=1/2*ae+ab+db+cb=...(do you still need the answer?)

OpenStudy (anonymous):

sklee, you probably ignore the important information that the container may not have a top!

OpenStudy (anonymous):

Ya, I just give it as included. So it depends on the full questions. If the container is open, then can just reduce (1/2)ae

OpenStudy (anonymous):

yes if you dont mind

OpenStudy (anonymous):

good!!!

OpenStudy (anonymous):

let's work out the answer as fast as possible

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so i add up all the sides which comes out to 7.9 right?

OpenStudy (anonymous):

6.9, probably.

OpenStudy (anonymous):

7.9,!! I'm sorry.

OpenStudy (anonymous):

ok, ya i got 7.9

OpenStudy (anonymous):

wait, wait, 0.5*2*1.5+(2+1.4+2)*1=6.9

OpenStudy (anonymous):

kinna dizzy...

OpenStudy (anonymous):

is 6.9 the answer? so we have to multiply by 0.5?

OpenStudy (anonymous):

yep.